Moving Proton in an Electric Field Part A A proton with kinetic energy of 1.02x105 eV is fired perpendicular to the face of a large plate that has a uniform charge density ofơ-+4.40 C/m2, what is the magnitude of the force on the proton? Submit Answer Tries 0/6 Part B How much work must the electric field do on the proton to bring it to rest? Submit Answer Tries 0/6 Part C From what distance should the proton be fired so that it stops right at the surface of the plate? Submit Answer Tries 0/6
Answer #1

Given,

KE = 1.02 x 10^5 eV ; sigma = 4.4 uC/m^2

We know that

E = sigma/2e0

E = 4.4 x 10^-6/(2 x 8.85 x 10^-12) = 2.5 x 10^5 N/C

F = q E

F = 1.6 x 10^-19 x 2.5 x 10^5 = 4 x 10^-14 N

hence, F = 4 x 10^-14 N

b)W = KE = 1.02 x 10^5 eV

W = 1.02 x 10^5 x 1.6 x 10^-19 = 1.632 x 10^-14 J

Hence, W = 1.632 x 10^-14 J

c)We know that

W = F x => x = F/W

x = (1.632 x 10^-14)/(4 x 10^-14) = 0.41 m

Hence, x = 0.41 m